A reaction has a rate constant of 0.0177 s-1 at 400.0 K. If the reaction has activation energy of 125 kJ/mol, calculate the rate constant at 500.0 K.A reaction has a rate constant of 0.0177 s-1 at 400.0 K. If the reaction has activation energy of 125 kJ/mol, calculate the rate constant at 500.0 K.
Answer #1

Answer -

Given,

Rate constant at 400.0 K (K1) = 0.0177 s-1

Activation Energy = 125 kJ/mol

Rate constant at 500.0 K (K2) = ?

We know that,

ln(K2/K1) = (Ea/R)[(1/T1)-(1/T2)]

where,

Ea = Activation Energy

K2 and K1 are rate constant

T1 and T2 are Temperatures

R = Rate Constant (0.008314 kJ/K mol)

Put the values,

ln(K2/0.0177 s-1) = (125 kJ/mol/0.008314 kJ/K mol)[(1/400.0 K)-(1/500.0 K)]

ln(K2/0.0177 s-1) = (15034.881 K) * (0.0005 K-1)

ln(K2/0.0177 s-1) = 7.517

K2/0.0177 s-1 = e7.517

K2/0.0177 s-1 = 1839.042

K2 = 1839.042 * 0.0177 s-1

K2 = 32.551 s-1 [ANSWER]